# v-representability and Hohenberg-Kohn theorem for non-interacting Schrödinger operators with distributional potentials in the one-dimensional torus

### v1

@@ 2.1. Notation
$\mathcal{D}_{N}\coloneqq\left\{\rho:\mathbb{T}\rightarrow\mathbb{R}\mbox{ such that}\quad\sqrt{\rho}\in\mathrm{H}^{1}(\mathbb{T}),\quad\int_{\mathbb{R}}\rho(x)\mathrm{d}x=N,\quad\mbox{and}\quad\rho>0\right\}$ (2.1)
@@
[…]
@@ 2.2. Main results
The second natural question that was left open in [SPR+24] is whether one can recover the potential $v$ from the ground-state density $\rho$. This question is not only natural but also specially relevant for (Kohn-Sham) DFT because it implies that the Kohn-Sham density-to-potential map $\rho\mapsto v^{\rm KS}(\rho)$ is single-valued. Moreover, as highlighted in [SPR+24, Corollary 19], the existence of a unique potential $v$ is also directly connected to the differentiability of the convex Lieb functional, which plays an important role in approximate schemes. Our next results provides an affirmative answer to this inverse problem for the case of non-interacting systems.
@@
In particular, as the set $\mathcal{D}_{1}^{\rm pure}(v)$ for any $v\in\mathcal{V}$ consists of a single density by Theorem 2.1, we obtain a complete solution to the pure-state $\mathcal{V}$-representability problem in the case of a single-particle ($N=1$). Moreover, in this case, the proof does not rely on the results from [SPR+24].
@@ 2.3. Outline of the paper
The proof of Theorem 2.1 consists of three steps and is carried out in Section 4. In the first step, we use classical results from the book by Reed and Simon [RS78] (cf. Theorems XVIII.43 and 45) to show that the ground-state of $h(v)$ is non-degenerate and strictly positive almost everywhere. In the second step, we apply Courant’s nodal domain theorem to conclude that the ground-state can not vanish in more than one point. If we were dealing with an interval $I$ with Dirichlet boundary conditions, these two steps would be enough to prove Theorem 2.1 because $I\setminus\{x_{0}\}$ consists of two connected components. In the torus, however, this is no longer true, and we need a third step. This last step uses a gluing argument to obtain a contradiction with Courant’s nodal domain theorem. More precisely, we use a gluing argument to construct ground-state densities that vanish on finitely many points, provided that a ground-state density vanishing on a single-point exists.
The proof of the Hohenberg-Kohn Theorem 2.5 is presented in Section 5. This proof consists of two main steps. The first step is the standard Hohenberg-Kohn argument, which shows that, if $H_{N}(v)$ and $H_{N}(v^{\prime})$ have the same ground-state density, they must have a mutual ground-state wave-function. After this step, the usual argument used in previous proofs of the Hohenberg-Kohn theorem [HK64, Lie83, PTC+23, Gar18] consists in dividing the difference of both Schrödinger equations by this mutual ground-state wave-function, which is possible in a (almost everywhere) pointwise sense by unique continuation results. In our setting, this is no longer possible because we are dealing with distributional potentials which are not pointwise defined and whose support may have Lebesgue measure zero. To overcome this difficulty, we appeal to the natural orbital decomposition of the ground-state. More precisely, we use the fact that, since the ground-state of the single-particle operator is non-degenerate, this state must be a natural orbital of any ground-state wave-function of the $N$-particles non-interacting system. Thus, we can reduce the situation to the case of a single-particle, where division by the ground-state is possible in an operator sense thanks to the strict positivity property in Theorem 2.1.
The proof of Theorem 2.8 is also presented in Section 5 and relies on the simple observation that any (real-valued) excited state $\varphi_{k}$ must vanish at a point, and therefore, adding a Dirac’s delta at that point still preserves $\varphi_{k}$ as an excited state.
@@ 3.1. Sobolev and dual spaces on the torus
First, let $\mathcal{B}(\mathrm{H}^{1}(T))$ denote the Banach space of bounded linear operators from $\mathrm{H}^{1}(\mathbb{T})$ to $\mathrm{H}^{1}(\mathbb{T})$, and note that, by Lemma 3.2, the map $M:\mathrm{H}^{1}(\mathbb{T})\rightarrow\mathcal{B}(\mathrm{H}^{1}(\mathbb{T}))$ given by
$\psi\in\mathrm{H}^{1}(\mathbb{T})\rightarrow M_{\psi}\in\mathcal{B}(\mathrm{H}^{1}(\mathbb{T})),\quad\mbox{where}\quad M_{\psi}(\varphi)=\psi\varphi$ (3.6)
@@
$\leq\lVert\ddot{g}\rVert_{\mathrm{L}^{\infty}(-C,C)}^{2}\lVert\delta\rVert_{\mathrm{H}^{1}}^{2}+\lVert g^{(3)}\rVert_{\mathrm{L}^{\infty}(-C,C)}^{2}\lVert\psi\rVert_{\mathrm{H}^{1}}^{2}\lVert\delta\rVert_{\mathrm{H}^{1}}^{2},$
@@ 3.2. Quadratic form definition of Hamiltonian
An useful consequence of this construction is that the quadratic form domain of $h(v)$ is the Sobolev space $\mathrm{H}^{1}(\mathbb{T})$. As this space is compactly embedded in $\mathrm{L}^{2}(\mathbb{T})$ by the standard Sobolev embedding theorem, the resolvent operator of $h(v)$ is a bounded operator form $\mathrm{L}^{2}(\mathbb{T})$ to $\mathrm{H}^{1}(\mathbb{T})$. Consequently, the resolvent is a compact operator and, by the spectral theorem, the spectrum of $h(v)$ is purely discrete, i.e., there exists a non-decresing sequence $\{\lambda_{k}\}_{k\in\mathbb{N}}$ and an orthonormal family $\{\varphi_{k}\}_{k\in N}$ such that
@@ 3.3. Single-particle density matrix and natural orbitals
where $a(\varphi)$ denotes the fermionic annihilation operator. It is well-known that the single-particle density matrix of any $N$-particles wave-function is a bounded, positive and trace-class operator satisfying
@@
Hence, $\gamma_{\Psi}$ is a compact self-adjoint operator on $\mathcal{H}_{1}$ and, by the spectral theorem, there exists an orthonormal family of eigenfunctions $\{\varphi_{j}\}_{j\leq M}$ and eigenvalues $\{\lambda_{k}\}_{k\leq M}$ such that
$0<\lambda_{k}\leq\lVert\Psi\rVert_{\mathcal{H}_{N}}^{2},\quad\sum_{k=1}^{M}\lambda_{k}=N\lVert\Psi\rVert_{\mathcal{H}_{N}}^{2},\quad\mbox{and}\quad\gamma_{\Psi}(x,y)=\sum_{k=1}^{M}\lambda_{k}\varphi_{k}(x)\overline{\varphi_{k}(y)}.$ (3.9)
Here $M\in\mathbb{N}\cup\{+\infty\}$, i.e., $M$ does not need to be finite. The eigenfunctions $\varphi_{k}$ are called the natural orbitals of $\Psi$, and $\lambda_{k}$ are called the associated (natural) occupation numbers.
@@
The proof follows from the fact that, since the spectrum of $h(v)$ is discrete (see previous subsection), there exists an orthonormal basis of eigenfunctions $\{\varphi_{j}\}_{j\in\mathbb{N}}$, and the Slater determinants
@@
where $a^{\ast}$ denotes the creation operator, form an eigenbasis of $H_{N}(v)$ for the $N$-particles space $\mathcal{H}_{N}$. ∎
@@ 4. Non-degeneracy of the ground-state
The first lemma shows that the ground-state of $h(v)$ is non-degenerate and almost everywhere strictly positive. Such results are sometimes called Perron–Frobenius theorem and can be understood as a unique continuation property but applies only to the ground-state of the single-particle system.
@@
To conclude, note that, since $\mathrm{L}^{2}(\mathbb{T})$ is dense [^5] Recall that the inclusion $\mathrm{L}^{2}(\mathbb{T})\subset\mathrm{H}^{-1}(\mathbb{T})$ is given by the Riesz mapping introduced in (3.3). in $\mathrm{H}^{-1}(\mathbb{T})$ (which follows from the spectral theorem) and $\mathrm{L}^{\infty}(\mathbb{T})$ is dense in $\mathrm{L}^{2}(\mathbb{T})$, there exists a sequence of bounded functions $v_{n}\in\mathrm{L}^{\infty}(\mathbb{T})$ such that
@@
Let $\varphi\in\mathrm{H}^{1}(\mathbb{T})$ be an eigenfunction of $h(v)$ with eigenvalue $\lambda$, then the number of nodal domains, i.e., the number of connected components of the (open) set $\{x\in\mathbb{T}:\varphi(x)\neq 0\}$ is less or equal than $n(\lambda)\coloneqq\sum_{\mu\leq\lambda}\dim\ker\left(h(v)-\mu\right)$.
@@
$q_{h(v)}(\varphi_{j},\varphi_{k})=\int_{\mathbb{T}}\overline{\nabla\varphi_{j}(x)}\nabla\varphi_{k}(x)\mathrm{d}x+v(\overline{\varphi_{j}}\psi_{k})=\int_{\mathbb{T}}\overline{\nabla\varphi_{j}(x)}\nabla\varphi_{k}(x)=0\quad\mbox{if $j\neq k$.}$
In particular, as $\varphi$ is an eigenfunction of $h(v)$ with eigenvalue $\lambda$ we have
@@
By Lemma 4.1 and Lemma 4.3 we see that $\varphi_{v}(x)$ can vanish on at most one point in $\mathbb{T}$. Indeed, if there were at least two points where $\varphi_{v}(x)=0$, there would be at least two connected components of $\{x\in\mathbb{T}:\varphi_{v}(x)>0\}$ which contradicts the non-degeneracy of the ground-state by Lemma 4.3. However, to complete the proof we need to show that $\varphi_{v}(x)$ can not vanish at a single point.
@@
So suppose that there exits a ground-state $\varphi_{v}\geq 0$ of $h(v)$ for some $v\in\mathbb{H}^{-1}(\mathbb{T})$ such that $\varphi_{v}(x)=0$ only at $x=0$. Without loss of generality, we also assume that the ground-state energy is 00. Then, we define
@@
$\begin{dcases}\widetilde{\varphi}(x)=\varphi_{v}(x)\quad&\mbox{for $|x\,\mathrm{mod}\,2\pi|\leq\pi/8$},\\ \widetilde{\varphi}(x)=\varphi_{v}(x-\pi\,\mathrm{mod}\,2\pi),\quad&\mbox{for $|x-\pi|\leq\pi/8$,}\\ \widetilde{\varphi}(x)>0,\quad\mbox{otherwise}.\end{dcases}$ (4.1)
@@
Next, let $\{p_{j}\}_{j=1}^{3}$ be a partition of the unity [^6] That such partition of unity exists is a well-known fact in differential geometry. In the simple case of a torus they can be constructed explicitly from any function $\eta$ as above. subordinate to the following cover of $\mathbb{T}$:
$I_{1}\coloneqq(-\pi/8,\pi/8)\,\mathrm{mod}\,2\pi,\quad I_{2}\coloneqq(7\pi/8,9\pi/8)\quad\mbox{and}\quad I_{3}\coloneqq\{x\in\mathbb{T}:|x\,\mathrm{mod}\,2\pi|\geq\pi/16,\quad|x-\pi|\geq\pi/16\}.$
In other words, each of the $p_{j}$ satisfies $p_{j}\in C^{\infty}_{c}(\mathbb{R})$, $p_{j}(x)=0$ for $x\neq I_{j}$, $p_{j}(x)\geq 0$, and together they satisfy
@@
$\widetilde{v}(\psi)$ $=v(p_{1}\psi)+v\left(\tau_{-\pi}(p_{2}\psi)\right)-\int_{\mathbb{T}}\nabla\widetilde{\varphi}(x)\nabla\left(\frac{p_{3}\psi}{\widetilde{\varphi}(x)}\right)(x)\mathrm{d}x$
$=v(p_{1}\psi)+v\left(\tau_{-\pi}(p_{2}\psi)\right)-\int_{I_{3}}\frac{\nabla\widetilde{\varphi}(x)}{\widetilde{\varphi}(x)}\left(\nabla p_{3}(x)\psi(x)+p_{3}(x)\nabla\psi(x)\right)-\frac{\nabla\widetilde{\varphi}(x)^{2}}{\widetilde{\varphi}(x)^{2}}p_{3}(x)\psi(x)\mathrm{d}x,$

### v2

@@ 2.1. Notation
$\mathcal{D}_{N}\coloneqq\left\{\rho:\mathbb{T}\rightarrow\mathbb{R}\mbox{ such that}\quad\sqrt{\rho}\in\mathrm{H}^{1}(\mathbb{T}),\quad\int_{\mathbb{T}}\rho(x)\mathrm{d}x=N,\quad\mbox{and}\quad\rho>0\right\}$ (2.1)
@@
$\mathcal{V}\coloneqq\{v\in\mathrm{H}^{-1}(\mathbb{T}):\quad v(f)\in\mathbb{R}\quad\mbox{for any real-valued function $f\in\mathrm{H}^{1}(\mathbb{T})$}\}.$ (2.2)
@@ 2.2. Main results
The second natural question that was left open in [SPR+24] is whether one can recover the potential $v$ from the ground-state density $\rho$. This question is not only natural but also specially relevant for (Kohn-Sham) DFT because it implies that the Kohn-Sham density-to-potential map $\rho\mapsto v^{\rm KS}(\rho)$ is single-valued. Moreover, as highlighted in [SPR+24, Corollary 19], the existence of a unique potential $v$ is also directly connected to the differentiability of the convex Lieb functional, which plays an important role in approximate schemes. Our next result provides an affirmative answer to this inverse problem for the case of non-interacting systems.
@@
In particular, as the set $\mathcal{D}_{1}^{\rm pure}(v)$ for any $v\in\mathcal{V}$ consists of a single density by Theorem 2.1, we obtain a complete solution to the pure-state $\mathcal{V}$-representability problem in the case of a single-particle ($N=1$). Moreover, the proof of this result does not rely on the results from [SPR+24].
@@ 2.3. Outline of the paper
The proof of Theorem 2.1 consists of three steps and is carried out in Section 4. In the first step, we use classical results from the book by Reed and Simon [RS78] (cf. Theorems XVIII.43 and 45) to show that the ground-state of $h(v)$ is non-degenerate and strictly positive almost everywhere. In the second step, we apply Courant’s nodal domain theorem to conclude that the ground-state can not vanish in more than one point. If we were dealing with an interval $I$ with Dirichlet boundary conditions, these two steps would be enough to prove Theorem 2.1 because $I\setminus\{x_{0}\}$ consists of two connected components. In the torus, however, this is no longer true and we need a third step. This last step uses a gluing argument to obtain a contradiction with Courant’s nodal domain theorem. More precisely, we use a gluing argument to construct ground-state densities that vanish on finitely many points, provided that a ground-state density vanishing on a single-point exists.
The proof of the Hohenberg-Kohn Theorem 2.5 is presented in Section 5. This proof consists of two main steps. The first step is the standard Hohenberg-Kohn argument, which shows that, if $H_{N}(v)$ and $H_{N}(v^{\prime})$ have the same ground-state density, they must have a mutual ground-state wave-function. After this step, the usual argument used in previous proofs of the Hohenberg-Kohn theorem [HK64, Lie83, PTC+23, Gar18] consists in dividing the Schrödinger equation
$\left(H_{N}(v)-H_{N}(v^{\prime})\right)\Psi=0$ (2.6)
by this mutual ground-state wave-function $\Psi$, which is possible in a (almost everywhere) pointwise sense by unique continuation results. In our setting, this is no longer possible [^3] In fact, in the single-particle case, the division by the ground-state argument is still possible (in an operator sense) thanks to the strict positivity of the ground-state (cf. Theorem 2.1) and the algebra property of $\mathrm{H}^{1}(\mathbb{T})$ (Lemma 3.2), see, e.g., the proof of Theorem 2.7. For $N\geq 2$, however, this argument does not apply as we have no control over the zero set of the $N$-particles ground-state $\Psi$. because we are dealing with distributional potentials which are not pointwise defined and whose support may have Lebesgue measure zero. To overcome this difficulty, we appeal to the fact that $h(v)$ has discrete spectrum, and combine a spectral representation of the dual space $\mathrm{H}^{-1}(\mathbb{T})$ (Lemma 3.5) with Theorem 2.1. More precisely, we show that, if $\Psi$ satisfies (2.6) in a weak sense, has finitely many natural orbitals, and one of these orbitals can be chosen strictly positive, then the difference of potentials $(v-v^{\prime})$ is much more regular than expected, namely, belongs to $\mathrm{H}^{1}(\mathbb{T})$. Combining this extra regularity with some linear algebra arguments, we can then show that $(v-v^{\prime})$ must be constant, which completes the proof of Theorem 2.5.
The proofs of Theorems 2.7 and 2.8 are also presented in Section 5. The former follows from the existence, strict positivity, and non-degeneracy of the ground-state, plus some fairly standard arguments to show smoothness of a map. The proof of the latter relies on the simple observation that any (real-valued) excited state $\varphi_{k}$ must vanish at a point, and therefore, adding a Dirac’s delta at that point still preserves $\varphi_{k}$ as an excited state.
@@ 3.1. Sobolev and dual spaces on the torus
Let $\mathcal{B}(\mathrm{H}^{1}(\mathbb{T}))$ denote the Banach space of bounded linear operators from $\mathrm{H}^{1}(\mathbb{T})$ to $\mathrm{H}^{1}(\mathbb{T})$. Then by Lemma 3.2, the map $M:\mathrm{H}^{1}(\mathbb{T})\rightarrow\mathcal{B}(\mathrm{H}^{1}(\mathbb{T}))$, given by
$\psi\in\mathrm{H}^{1}(\mathbb{T})\rightarrow M_{\psi}\in\mathcal{B}(\mathrm{H}^{1}(\mathbb{T}))\quad\mbox{where}\quad M_{\psi}(\varphi)=\psi\varphi,$ (3.6)
@@
$\leq\lVert\ddot{g}\rVert_{\mathrm{L}^{\infty}(-C,C)}^{2}\lVert\delta\rVert_{\mathrm{H}^{1}}^{4}+\lVert g^{(3)}\rVert_{\mathrm{L}^{\infty}(-C,C)}^{2}\lVert\psi\rVert_{\mathrm{H}^{1}}^{2}\lVert\delta\rVert_{\mathrm{H}^{1}}^{4},$
@@ 3.2. Quadratic form definition of Hamiltonian
An useful consequence of this construction is that the quadratic form domain of $h(v)$ is the Sobolev space $\mathrm{H}^{1}(\mathbb{T})$. As this space is compactly embedded in $\mathrm{L}^{2}(\mathbb{T})$ by the standard Sobolev embedding theorem, the resolvent operator of $h(v)$ is a bounded operator form $\mathrm{L}^{2}(\mathbb{T})$ to $\mathrm{H}^{1}(\mathbb{T})$. Consequently, the resolvent is a compact operator and, by the spectral theorem, the spectrum of $h(v)$ is purely discrete, i.e., there exists a non-decresing sequence $\{\lambda_{k}\}_{k\in\mathbb{N}}$ and an $\mathrm{L}^{2}$-orthonormal basis $\{\varphi_{k}\}_{k\in\mathbb{N}}$ such that
@@
The above spectral decomposition of $h(v)$ will be useful to prove the Hohenberg-Kohn theorem 2.5 in Section 5. To be more precise, there we shall use the following representation of distributions in $\mathrm{H}^{-1}(\mathbb{T})$.
Lemma 3.5 (Spectral representation of dual Sobolev space).
Let $v\in\mathrm{H}^{-1}(\mathbb{T})$ and $\{\varphi_{k}\}_{k\in\mathbb{N}}$ be an orthonormal basis of eigenfunctions of $h(v)$ with corresponding eigenvalues $\{\lambda_{k}\}_{k\geq 1}$. Then a distribution $f\in\mathcal{D}^{\prime}(\mathbb{T})$ belongs to $\mathrm{H}^{-1}(\mathbb{T})$ if and only if there exists a sequence $\{c_{j}\}_{j\in\mathbb{N}}$ such that
$\sum_{j\in\mathbb{N}}(1-\lambda_{1}+\lambda_{j})^{-1}|c_{j}|^{2}<\infty\quad\mbox{and}\quad f(\varphi)=\sum_{j=1}^{\infty}c_{j}\langle\varphi_{j},\varphi\rangle,\quad\mbox{for any $\varphi\in\mathrm{H}^{1}(\mathbb{T})$.}$ (3.9)
Moreover, in this case $c_{j}=f(\varphi_{j})$.
Proof.
The proof is rather standard, but we sketch the arguments for convenience of the reader. First, note that any $\varphi\in\mathrm{L}^{2}(\mathbb{T})$, can be written as $\varphi=\sum_{j\geq 1}a_{j}\varphi_{j}$ where $a_{j}\coloneqq\langle\varphi_{j},\varphi\rangle$. Since $h(v)$ has quadratic form domain $\mathrm{H}^{1}(\mathbb{T})$ (see also (3.8)), we find that
$\lVert\varphi\rVert_{\mathrm{H}^{1}(\mathbb{T})}^{2}\sim\langle\varphi,\left(h(v)-\lambda_{1}+1\right)\varphi\rangle=\sum_{j=1}^{\infty}|a_{j}|^{2}(\lambda_{j}-\lambda_{1}+1),$
where $\sim$ denotes equivalence of norms. Hence, $\varphi\in\mathrm{H}^{1}(\mathbb{T})$ if and only if $\sum|a_{j}|^{2}(1-\lambda_{1}+\lambda_{j})<\infty$. Therefore, any $f$ of the form (3.9) belongs to $\mathrm{H}^{-1}(\mathbb{T})$ by Cauchy-Schwarz, i.e.,
$f(\varphi)=\sum_{j=1}^{\infty}c_{j}a_{j}\leq\left(\sum_{j=1}^{\infty}\frac{|c_{j}|^{2}}{(1+\lambda_{j}-\lambda_{1})}\right)^{\frac{1}{2}}\left(\sum_{j\geq 1}|a_{j}|^{2}(1+\lambda_{j}-\lambda_{1})\right)^{\frac{1}{2}}\lesssim\lVert\varphi\rVert_{\mathrm{H}^{1}(\mathbb{T})}.$
Conversely, if $f\in\mathrm{H}^{-1}(\mathbb{T})$, then we can define $c_{j}\coloneqq f(\varphi_{j})$, and conclude by using the simple fact that
$\left|\sum_{j=1}^{\infty}c_{j}a_{j}\right|\lesssim\left(\sum_{j=1}^{\infty}|a_{j}|^{2}(1+\lambda_{j}-\lambda_{1})\right)^{\frac{1}{2}},\quad\mbox{for any sequence $\{a_{j}\}_{j\in\mathbb{N}}$,}$
if and only if $\{c_{j}\}_{j\in\mathbb{N}}$ satisfies (3.9). ∎
@@ 3.3. Single-particle density matrix and natural orbitals
where $a(\varphi)$ denotes the usual fermionic annihilation operator,
$a(\varphi):\mathcal{H}_{N}\rightarrow\mathcal{H}_{N-1},\quad\left(a(\varphi)\Psi\right)(x_{1},...,x_{N-1})=\sqrt{N}\int_{\mathbb{T}}\overline{\varphi(x)}\Psi(x,x_{1},...,x_{N-1})\mathrm{d}x.$ (3.10)
It is well-known that the single-particle density matrix of any $N$-particles wave-function is a bounded, positive and trace-class operator satisfying
@@
Hence, $\gamma_{\Psi}$ is a compact self-adjoint operator on $\mathcal{H}_{1}$ and, by the spectral theorem, there exists an orthonormal family of eigenfunctions $\{\varphi_{j}\}_{j\leq M}$ and eigenvalues $\{n_{k}\}_{k\leq M}$ such that
$0<n_{k}\leq\lVert\Psi\rVert_{\mathcal{H}_{N}}^{2},\quad\sum_{k=1}^{M}n_{k}=N\lVert\Psi\rVert_{\mathcal{H}_{N}}^{2},\quad\mbox{and}\quad\gamma_{\Psi}(x,y)=\sum_{k=1}^{M}n_{k}\overline{\varphi_{k}(x)}\varphi_{k}(y).$ (3.11)
Here $M\in\mathbb{N}\cup\{+\infty\}$, i.e., $M$ does not need to be finite. The eigenfunctions $\varphi_{k}$ are called the natural orbitals of $\Psi$, and $n_{k}$ are called the associated (natural) occupation numbers.
Remark 3.7 (Non-uniqueness of natural orbitals).
Note that the natural orbitals are not uniquely defined, as any linear combination of the orbitals with the same occupation number is again a natural orbital.
For our proofs, we shall need two properties of the natural orbitals. The first one is the general fact that the natural orbitals of eigenfunctions of non-interacting operators can be chosen as eigenfunctions of the single-particle operator.
@@
The proof follows from the fact that, since the spectrum of $h(v)$ is discrete (see previous subsection), there exists an orthonormal basis of eigenfunctions $\{\varphi_{j}\}_{j\in\mathbb{N}}$, and therefore, the Slater determinants
@@
where $1\in\mathbb{C}=\mathcal{H}_{0}$ is the vacuum state and $a(\varphi)^{\ast}:\mathcal{H}_{N-1}\rightarrow\mathcal{H}_{N}$ is the creation operator
$\left(a(\varphi)^{\ast}\Psi\right)(x_{1},...,x_{N})=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}(-1)^{1+i}\varphi(x_{i})\Psi(x_{1},...,x_{i-1},x_{i+1},...,x_{N}),$ (3.12)
form an eigenbasis of $H_{N}(v)$ for the $N$-particles space $\mathcal{H}_{N}$. ∎
The second result we shall need is the following simple formula for the overlapping density of a wave-function with its single-excitations.
Lemma 3.9 (Overlapping density with single-excitations).
Let $\Psi\in\mathcal{H}_{N}$ and $\{\varphi_{j}\}_{j=1}^{M}$ be its natural orbitals. Then for $\Phi=a(\psi)^{\ast}a(\varphi_{k})\Psi$, where $\psi\perp\{\varphi_{j}\}_{j\leq M}$ and $a(\psi)^{\ast}$ and $a(\varphi_{k})$ are the creation and annihilation operators defined in (3.12) and (3.10), we have
$\rho_{\Psi,\Phi}(x)=\int_{\mathbb{T}^{N-1}}\overline{\Psi(x,x_{2},...,x_{N})}\Phi(x,x_{2},...,x_{N})\mathrm{d}x_{2}...\mathrm{d}x_{N}=n_{k}\overline{\varphi_{k}(x)}\psi(x),$ (3.13)
where $n_{k}>0$ is the occupation number of $\varphi_{k}$.
Proof.
First note that since $\psi\perp\{\varphi_{j}:j\leq M\}$ can be assumed to be normalized, we can extend $\{\varphi_{j}\}_{j\leq M}$ to an orthornomal basis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ such that $\varphi_{M+1}=\psi$. Now consider $v\in\mathrm{L}^{\infty}(\mathbb{T})$ and denote by $V_{N}$ the $N$-particle operator of multiplication by
$V_{N}(x_{1},...,x_{N})=\sum_{j=1}^{N}v(x_{j}).$
Then from the second quantization representation
$V_{N}=\sum_{i,j\geq 1}v_{ij}a_{i}^{\ast}a_{j},\quad\mbox{where $a_{i}=a(\varphi_{i})$, $a_{j}^{\ast}=a(\varphi_{j})^{\ast}$, and $v_{ij}=\langle\varphi_{i},v\varphi_{j}\rangle$},$
and the canonical anti-commutation relations (CAR)
$a_{j}a_{k}=-a_{k}a_{j},\quad\mbox{and}\quad a_{k}a_{j}^{\ast}=\delta_{kj}-a_{j}^{\ast}a_{k},\quad\mbox{where $\delta_{ij}$ is the Kronecker delta,}$,
we find that
$\langle\Psi,V_{N}a(\psi)^{\ast}a(\varphi_{k})\Psi\rangle=\sum_{i,j}v_{ij}\langle\Psi,a_{i}^{\ast}a_{j}a_{M+1}^{\ast}a_{k}\Psi\rangle=\sum_{i}v_{iM+1}\langle a_{i}\Psi,a_{k}\Psi\rangle+\sum_{i,j}v_{ij}\langle a_{i}a_{M+1}\Psi,a_{j}a_{k}\Psi\rangle.$
We can now use that $a_{k}\Psi=0$ for any $k>M$ (since $\varphi_{k}\perp\{\varphi_{j}\}_{j\leq M}$ which are the natural orbitals of $\Psi$) to obtain
$\langle\Psi,V_{N}\Phi\rangle=\sum_{i=1}^{M}v_{iM+1}\langle\varphi_{i},\gamma_{\Psi}\varphi_{k}\rangle=n_{k}v_{kM+1}=\int_{\mathbb{T}}v(x)n_{k}\overline{\varphi}_{k}(x)\psi(x)\mathrm{d}x.$ (3.14)
On the other hand, we have
$\langle\Psi,V_{N}\Phi\rangle=\int_{\mathbb{T}}v(x)\rho_{\Psi,\Phi}(x)\mathrm{d}x.$ (3.15)
As (3.14) and (3.15) holds for any $v\in\mathrm{L}^{\infty}(\mathbb{T})$, we obtain (3.13). ∎
@@ 4. Non-degeneracy of the ground-state
The first lemma shows that the ground-state of $h(v)$ is non-degenerate and almost everywhere strictly positive. This kind of result is sometimes called Perron–Frobenius theorem; it can be understood as a unique continuation property but applies only to the ground-state of the single-particle system.
@@
To conclude, note that, since $\mathrm{L}^{2}(\mathbb{T})$ is dense [^6] Recall that the inclusion $\mathrm{L}^{2}(\mathbb{T})\subset\mathrm{H}^{-1}(\mathbb{T})$ is given by the Riesz mapping introduced in (3.3). in $\mathrm{H}^{-1}(\mathbb{T})$ (by Lemma 3.5) and $\mathrm{L}^{\infty}(\mathbb{T})$ is dense in $\mathrm{L}^{2}(\mathbb{T})$, there exists a sequence of bounded functions $v_{n}\in\mathrm{L}^{\infty}(\mathbb{T})$ such that
@@
Let $\varphi\in\mathrm{H}^{1}(\mathbb{T})$ be an eigenfunction of $h(v)$ with eigenvalue $\lambda$, then the number of nodal domains of $\varphi$, i.e., the number of connected components of the (open) set $\{x\in\mathbb{T}:\varphi(x)\neq 0\}$ is less or equal than $n(\lambda)\coloneqq\sum_{\mu\leq\lambda}\dim\ker\left(h(v)-\mu\right)$.
@@
$q_{h(v)}(\varphi_{j},\varphi_{k})=\int_{\mathbb{T}}\overline{\nabla\varphi_{j}(x)}\nabla\varphi_{k}(x)\mathrm{d}x+v(\overline{\varphi_{j}}\varphi_{k})=\int_{\mathbb{T}}\overline{\nabla\varphi_{j}(x)}\nabla\varphi_{k}(x)=0\quad\mbox{if $j\neq k$.}$
In particular, as $\varphi=\sum\varphi_{k}$ is an eigenfunction of $h(v)$ with eigenvalue $\lambda$ we have
@@
By Lemma 4.1 and Lemma 4.3, we see that $\varphi_{v}(x)$ can vanish on at most one point in $\mathbb{T}$. Indeed, if there were at least two points where $\varphi_{v}(x)=0$, there would be at least two connected components of $\{x\in\mathbb{T}:\varphi_{v}(x)>0\}$ which contradicts the non-degeneracy of the ground-state by Lemma 4.3. However, to complete the proof we need to show that $\varphi_{v}(x)$ can not vanish at a single point.
@@
So suppose that there exists a ground-state $\varphi_{v}\geq 0$ of $h(v)$ for some $v\in\mathbb{H}^{-1}(\mathbb{T})$ such that $\varphi_{v}(x)=0$ only at $x=0$. Without loss of generality, we also assume that the ground-state energy is 00. Then, we define
@@
$\begin{dcases}\widetilde{\varphi}(x)=\varphi_{v}(x)\quad&\mbox{for $|x\,\mathrm{mod}\,2\pi|\leq\pi/8$},\\ \widetilde{\varphi}(x)=\varphi_{v}(x-\pi),\quad&\mbox{for $|x-\pi|\leq\pi/8$,}\\ \widetilde{\varphi}(x)>0,\quad\mbox{otherwise}.\end{dcases}$ (4.1)
@@
Next, let $\{p_{j}\}_{j=1}^{3}$ be a partition of the unity [^7] That such partition of unity exists is a well-known fact in differential geometry. In the simple case of a torus they can be constructed explicitly from any function $\eta$ as above. subordinate to the following open cover of $\mathbb{T}$:
$I_{1}\coloneqq(-\pi/8,\pi/8)\,\mathrm{mod}\,2\pi,\quad I_{2}\coloneqq(7\pi/8,9\pi/8)\quad\mbox{and}\quad I_{3}\coloneqq\{x\in\mathbb{T}:|x\,\mathrm{mod}\,2\pi|>\pi/16,\quad|x-\pi|>\pi/16\}.$
In other words, each of the $p_{j}$ satisfies $p_{j}\in C^{\infty}(\mathbb{T})$, $p_{j}(x)=0$ for $x\not\in I_{j}$, $p_{j}(x)\geq 0$, and together they satisfy
@@
$\widetilde{v}(\psi)$ $=v(p_{1}\psi)+v\left(\tau_{\pi}(p_{2}\psi)\right)-\int_{\mathbb{T}}\nabla\widetilde{\varphi}(x)\nabla\left(\frac{p_{3}\psi}{\widetilde{\varphi}}\right)(x)\mathrm{d}x$
$=v(p_{1}\psi)+v\left(\tau_{\pi}(p_{2}\psi)\right)-\int_{I_{3}}\frac{\nabla\widetilde{\varphi}(x)}{\widetilde{\varphi}(x)}\left(\nabla p_{3}(x)\psi(x)+p_{3}(x)\nabla\psi(x)\right)-\frac{\nabla\widetilde{\varphi}(x)^{2}}{\widetilde{\varphi}(x)^{2}}p_{3}(x)\psi(x)\mathrm{d}x,$